# 2.5cm2.5cm2.5cm.8cm

### v1

@@ Abstract
In policy debates concerning the governance and regulation of Artificial Intelligence (AI), both the Precautionary Principle (PP) and the Innovation Principle (IP) are advocated by their respective interest groups. Do these principles offer wholly incompatible and contradictory guidance? Does one necessarily negate the other? I argue here that provided attention is restricted to weak-form PP and IP, the answer to both of these questions is “No.” The essence of these weak formulations is the requirement to fully account for type-I error costs arising from erroneously preventing the innovation’s diffusion through society (i.e. mistaken regulatory red-lighting) as well as the type-II error costs arising from erroneously allowing the innovation to diffuse through society (i.e. mistaken regulatory green-lighting). Within the Signal Detection Theory (SDT) model developed here, weak-PP red-light (weak-IP green-light) determinations are optimal for sufficiently small (large) ratios of expected type-I to type-II error costs. For intermediate expected cost ratios, an amber-light ‘wait-and-monitor’ policy is optimal. Regulatory sandbox instruments allow AI testing and experimentation to take place within a structured environment of limited duration and societal scale, whereby the expected cost ratio falls within the ‘wait-and-monitor’ range. Through sandboxing regulators and innovating firms learn more about the expected cost ratio, and what respective adaptations — of regulation, of technical solution, of business model, or combination thereof, if any — are needed to keep the ratio out of the weak-PP red-light zone.
@@ 1 INTRODUCTION
The key novel insight underpinning this paper is that one of the two corner solutions in Signal Detection Theory (SDT) corresponds to non-postponement of weak-PP red-light intervention, while the remaining corner solution corresponds to non-postponement of weak-IP green-lighting. The corner solutions occur when the expected error cost structure is highly asymmetric. This paper addresses the question of just how asymmetric the cost structure must be in order to trigger weak-PP red-light (or weak-IP green-light) intervention. The answer depends on the distinguishing power of current AI-risk science, captured in the discriminability parameter $d^{\prime}$, as well as an inferential-surprise-tolerance parameter $\epsilon$ [^3] For details, see Section 4, p 4. Superficial resemblance aside, the latter parameter should not be confused with a de minismis threshold.
Careful consideration of SDT mathematical structure yields further insights into the consequences of interactions between base rates, expected error cost structure, and discriminability $d^{\prime}$. For instance when public trust in science is eroded, from the public’s standpoint the relevant discriminability parameter is zero. If one sets $d^{\prime}=0$ in SDT, it follows that for all less-than-unity ($<1$) expected error cost ratios, weak-PP red-light intervention is optimal, whereas for all greater-than-unity ($>1$) expected error cost ratios, weak-IP green-lighting is optimal. And if the erosion of public trust extends to the work of risk professionals and regulatory institutions, then each respective interest group will prioritize their own perception of the expected error cost ratio and the associated red-light or green-light determination — i.e. the result is societal polarization by interest group, even without the effect of non-comparable protected values.
@@ 3 Optimal inferential thresholds for policy decisions
Actual $H_{1}$    $TP$    $FN$
$H_{0}$ $FP$ $TN$
@@ 4 Non-postponement
Figure 3: Iso-expected cost line (in red) that is tangent with the ROC curve (in cyan) within the $\delta_{1}=0.004$ neighborhood of (0,0) (in magenta).
@@ 6.1 Implications for the PP
Just because an industrial innovation poses a credible risk of serious damage to the environment, or an AI innovation poses a credible risk to current technical and economic structures, it does not necessarily follow that the optimal operating point will fall within the $\delta_{1}$-neighborhood of (1,1). It is entirely possible for there to be serious risks of damage, yet when expected incremental misclassification costs are computed according to equation (3.5), the resulting optimal operating point remains within the amber-light ‘wait-and-monitor’ interior of the ROC space. For smaller discriminabilities and AUCs, however, the likelihood of the optimal operating point reaching the neighbourhood of a corner solution increases, ceteris paribus. In the limit, as $d^{\prime}\to 0$ and AUC0.5absent0.5$\to$ 0.5→ 0.5, even minute asymmetry between the expected costs of false positives and false negatives will lead to a corner solution. Indeed when societal trust in the ability of science to deliver credible, politically even-handed positive discriminability $d^{\prime}>0$ is eroded, then even under the SDT-based model one expects to see polarization. Part of society will have steeper expected misclassification cost ratios, a (0,0) corner solution, and will argue for the removal of all restrictions on AI innovation. Another part of society will have flatter expected misclassification cost ratios, a (1,1) corner solution, and will argue for the shuttering of AI innovation.
@@ 7 CONCLUSION
[…]
@@ A.1 Observations
Fisher introduced significance testing and the concept of a $p$-value, i.e. the probability that a test-statistic $T=t(X)$,313131computed on observed data drawn from a continuous distribution $X\sim f(x|\theta)$ on support $\mathbb{R}$ equals or exceeds the observed value $t(x)$ given that the null hypothesis $H_{0}:\theta=\theta_{0}$ is true, i.e. $p=P(t(X)\geq t(x)|H_{0})$. In Fisher’s approach to significance testing, there is no explicit alternative hypothesis under consideration. This is because there are innumerable different conceivable alternative hypotheses. Fisher views the alternative hypothesis — and therefore any quantities derived from it, such as statistical power — as ‘unknown’. Although Fisher believed that $p$-values require researchers’ subjective interpretation, his early expositions advocated using $p<0.05$ (i.e. a 5% significance level) as the standard for concluding that there is evidence against $H_{0}$.
Fisher (1925): The value for which P $=$ .05, or 1 in 20, is 1.96 or nearly 2; it is convenient to take this point as a limit in judging whether a deviation is to be considered significant or not. … …We shall not often be astray if we draw a conventional line at 0.05 … .
@@
It is customary therefore to assign a bound to the probability of incorrectly rejecting [$H_{0}$] when it is true and to attempt to minimize the other probability subject to this condition. … …The choice of a level of significance $\alpha$ is usually somewhat arbitrary… …Standard values, such as .01 or .05, were originally chosen to effect a reduction in the tables needed for carrying out various test [sic]. By habit, and because of the convenience of standardization in providing a common frame of reference, these values gradually became entrenched as the conventional levels to use. (Lehmann and Romano, 2005)

### v2

@@ Abstract
In policy debates concerning the governance and regulation of Artificial Intelligence (AI), both the Precautionary Principle (PP) and the Innovation Principle (IP) are advocated by their respective interest groups. Do these principles offer wholly incompatible and contradictory guidance? Does one necessarily negate the other? I argue here that provided attention is restricted to weak-form PP and IP, the answer to both of these questions is “No.” The essence of these weak formulations is the requirement to fully account for type-I error costs arising from erroneously preventing the innovation’s diffusion through society (i.e. mistaken regulatory red-lighting) as well as the type-II error costs arising from erroneously allowing the innovation to diffuse through society (i.e. mistaken regulatory green-lighting). Within the Signal Detection Theory (SDT) model developed here, weak-PP red-light (weak-IP green-light) determinations are optimal for sufficiently small (large) ratios of expected type-I to type-II error costs. For intermediate expected cost ratios, an amber-light ‘wait-and-monitor’ policy is optimal. Regulatory sandbox instruments allow AI testing and experimentation to take place within a structured environment of limited duration and societal scale, whereby the expected cost ratio falls within the ‘wait-and-monitor’ range. Through sandboxing regulators and innovating firms learn more about the expected cost ratio, and what respective adaptations — of regulation, of technical solution, of business model, or combination thereof, if any — are needed to keep the ratio out of the weak-PP red-light zone. Nevertheless AI foundation models are ill-suited for regulatory sandboxing as their general-purpose nature precludes credible identification of misclassification costs.
@@ 1 INTRODUCTION
The key novel insight underpinning this paper is that one of the two corner solutions in Signal Detection Theory (SDT) corresponds to non-postponement of weak-PP red-light intervention, while the remaining corner solution corresponds to non-postponement of weak-IP green-lighting. The corner solutions occur when the expected error cost structure is highly asymmetric. This paper addresses the question of just how asymmetric the cost structure must be in order to trigger weak-PP red-light (or weak-IP green-light) intervention. The answer depends on the distinguishing power of current AI-risk science, captured in the discriminability parameter $d^{\prime}$, as well as an inferential-surprise-tolerance parameter $\epsilon$. [^3] For details, see Section 4, p 4. Superficial resemblance aside, the latter parameter should not be confused with a de minismis threshold.
Careful consideration of SDT mathematical structure yields further insights into the consequences of interactions between base rates, expected error cost structure, and discriminability $d^{\prime}$. For instance when public trust in science is eroded, from the public’s standpoint the relevant discriminability parameter is zero. If one sets $d^{\prime}=0$ in SDT, it follows that for all less-than-unity (<1<1) expected error cost ratios, weak-PP red-light intervention is optimal, whereas for all greater-than-unity (>1>1) expected error cost ratios, weak-IP green-lighting is optimal. And if the erosion of public trust extends to the work of risk professionals and regulatory institutions, then each respective interest group will prioritize their own perception of the expected error cost ratio and the associated red-light or green-light determination — i.e. the result is societal polarization by interest group, even without the effect of non-comparable protected values.
@@ 3 Optimal inferential thresholds for policy decisions
Actual $H_{1}$    TPTP    FNFN
$H_{0}$ FPFP TNTN
@@ 4 Non-postponement
Figure 3: Iso-expected cost line (in red) that is tangent with the ROC curve (in cyan) within the $\delta_{0}=0.004$ neighborhood of (0,0) (in magenta).
@@ 6.1 Implications for the PP
[…]
@@ 7 CONCLUSION
The first of these, which can be interpreted as the reason for the existence of the PP, is the bridging of the disparate inferential conventions and requirements of science (NHST) and policy decision making. NHST inferential practices followed in science are intended to hold type-I error at the fixed, conventional level of $\alpha=.05$. However policy decision making requires an inferential threshold that reflects operative cost tradeoffs. The weak PP, understood as a verbal formulation of an SDT corner solution, serves as a patch for this disparity between the inferential conventions of science and the needs of practical policy decision making.
@@ A.1 Observations
Fisher introduced significance testing and the concept of a $p$-value, i.e. the probability that a test-statistic T=t(X)T=t(X),313131computed on observed data drawn from a continuous distribution $X\sim f(x|\theta)$ on support $\mathbb{R}$ equals or exceeds the observed value $t(x)$ given that the null hypothesis $H_{0}:\theta=\theta_{0}$ is true, i.e. $p=P(t(X)\geq t(x)|H_{0})$. In Fisher’s approach to significance testing, there is no explicit alternative hypothesis under consideration. This is because there are innumerable different conceivable alternative hypotheses. Fisher views the alternative hypothesis — and therefore any quantities derived from it, such as statistical power — as ‘unknown’. Although Fisher believed that $p$-values require researchers’ subjective interpretation, his early expositions advocated using $p<0.05$ (i.e. a 5% significance level) as the standard for concluding that there is evidence against $H_{0}$.
Fisher (1925): The value for which P == .05, or 1 in 20, is 1.96 or nearly 2; it is convenient to take this point as a limit in judging whether a deviation is to be considered significant or not. … …We shall not often be astray if we draw a conventional line at 0.05 … .
@@
It is customary therefore to assign a bound to the probability of incorrectly rejecting $H_{0}$] when it is true and to attempt to minimize the other probability subject to this condition. … …The choice of a level of significance $\alpha$ is usually somewhat arbitrary… …Standard values, such as .01 or .05, were originally chosen to effect a reduction in the tables needed for carrying out various test [sic]. By habit, and because of the convenience of standardization in providing a common frame of reference, these values gradually became entrenched as the conventional levels to use. (Lehmann and Romano, 2005)